OpenStudy (anonymous):

is -7 + 9 = -9 + 7 true, false, or open?

OpenStudy (teddyiswatshecallsme):

I don't understand what it is asking......

OpenStudy (anonymous):

well thats what it says on my quiz so i have no other way of asking it.

OpenStudy (teddyiswatshecallsme):

Ok but what is it talking about false,true, or open?

OpenStudy (anonymous):

the course is over introduction to equations.

OpenStudy (anonymous):

I hope you guys are joking. -7+9=-9+7 2 =/= -2 false.